# Integrating Deep-Learning-Based Magnetic Model and Non-Collinear Spin-Constrained Method: Methodology, Implementation and Application

### v2

@@ Abstract
We propose a non-collinear spin-constrained method that generates training data for deep-learning-based magnetic model, which provides a powerful tool for studying complex magnetic phenomena that requires large-scale simulations at the atomic level. First, we propose a projection method for calculating atomic magnetic moments by applying a radial truncation to numerical atomic orbitals. We then implement a Lagrange multiplier method that yields the magnetic torques of atoms by constraining the magnitude and direction of atomic magnetic moments. The method is implemented in ABACUS with both plane wave basis and numerical atomic orbital basis. We benchmark the iron (Fe) systems and analyze differences from calculations with the plane wave basis and numerical atomic orbitals basis in describing magnetic energy barriers. Based on more than 30,000 first-principles data with the information of magnetic torque, we train a deep-learning-based magnetic model DeePSPIN for the Fe system. By utilizing the model in large-scale molecular dynamics simulations, we successfully predict Curie temperatures of $\alpha$-Fe close to experimental values.
I Introduction
@@
The study of magnetic materials has been a cornerstone of condensed matter physics, with implications ranging from fundamental science to technological applications. Density Functional Theory (DFT) Hohenberg and Kohn (1964); Kohn and Sham (1965) has emerged as a powerful tool for understanding the electronic and magnetic properties of these materials Szilva et al. (2023). However, accurate description of excited magnetic states within DFT has remained a formidable challenge due to the complex interplay between electron and lattice behaviors.
@@
The development of constrained Density Functional Theory (cDFT) has marked a significant advancement in this field  Dederichs et al. (1984a); Kaduk et al. (2012). In the pioneering work of Dederichs et al. in 1984 Dederichs et al. (1984a), DFT was extended to arbitrary constraints through the introduction of the Lagrange multiplier method. Since then, several cDFT methods have been used to study excited states of charge distribution Dederichs et al. (1984a); Wu and Van Voorhis (2005) or magnetization Stocks et al. (1998); Kurz et al. (2004); Cuadrado et al. (2018); Gebauer and Baroni (2000); Ma and Dudarev (2015); Hegde et al. (2020a); Cai et al. (2023). Introducing penalty functions is commonly adopted to implement any constraint. Gebauer and Baroni (2000); Ma and Dudarev (2015). Wu and Van Voorhis (2005) Wu and Van Voorhis (2005) introduced an efficient algorithm, which allows for the use of a double-loop method to find the effective Lagrange multiplier that satisfies the constraints. In addition, the cDFT methods have been implemented based on different basis sets, such as the real-space grid Gebauer and Baroni (2000), the full-potential linearized augmented plane-wave (FLAPW) Kurz et al. (2004), the projector augmented-wave (PAW) Ma and Dudarev (2015); Hegde et al. (2020a); Cai et al. (2023), and the numerical atomic orbitals (NAOs) Cuadrado et al. (2018). These efforts provide powerful tools in understanding charge and magnetization fluctuations in solids, predicting spin-dependent phenomena, and characterizing electron transfer reactions in molecules Kaduk et al. (2012); Stocks et al. (1998); Kurz et al. (2004); Wu and Van Voorhis (2005); Ma and Dudarev (2015); Cuadrado et al. (2018); Hegde et al. (2020a); Cai et al. (2023).
Reviewing the numerous spin-cDFT methods, one of main differences lies in the definitions of atomic magnetic moments. One category is the Mulliken population method Cuadrado et al. (2018), which, while maintaining the sum rule, may introduce contributions from non-local orbitals. Another approach is based on the local orbital projection Hegde et al. (2020a); Cai et al. (2023), which preserves better the locality of magnetic moments. In this work, we implement a spin-constrained method with both plane-wave (PW) basis and numerical atomic orbital basis within the DFT package ABACUS Chen et al. (2010); Li et al. (2016); Lin et al. (2020); Zhou et al. (2025). Benefiting from the locality of NAOs, magnetic moments are obtained through the projection of atomic orbitals. By appropriately truncating the local atomic orbitals and smoothing near the truncation, we find that the modulated atomic orbitals provide satisfactory performance in stability and convergence during cDFT calculations. The spin-cDFT is fully optimized using the double-loop approach Wu and Van Voorhis (2005). Importantly, the optimized Lagrange multipliers can serve as magnetic torques for atoms.
@@
In recent years, the integration of deep learning models with DFT has offered unprecedented opportunities for transferring first-principles accuracy to larger scales Zhang et al. (2024, 2023); Qi et al. (2024); Chen and Ong (2022); Deng et al. (2023); Batatia et al. (2023); Choudhary et al. (2023); Choudhary and DeCost (2021). The available databases Jain et al. (2013); Talirz et al. (2020); ato ; Barroso-Luque et al. (2024) contain a large amount of DFT data that can be used to train machine learning models for different purposes. Liu et al. (2024); Wu et al. (2023a, b). Recently, some magnetic models Rinaldi et al. (2024); Yang et al. (2024) have emerged that incorporate the degrees of freedom of spin into the existing neural networks and can perform large-scale magnetic dynamics simulations. However, these models suffer from a lack of data since most existing DFT data in datasets are non-magnetic or only include ground-state magnetic results. Ground-state data cannot meet the sampling requirements of magnetic models with an additional degree of freedom spin. From this perspective, spin-cDFT methods not only act as powerful first-principles tools for studying magnetic excited states at the atomic scale but also provide training data for machine-learning-based magnetic models. Additionally, incorporating magnetic torque is advantageous for magnetic models that can describe magnetic field interactions. This requires that the analytical form of the magnetic torque can be satisfied through the definition of the constraints. However, most existing methods of cDFT, especially those with penalty functions, do not meet this requirement. In this work, we implement a spin-cDFT method within the ABACUS package, which provides energies, atomic forces, stresses, and magnetic torques for any magnetic excited state, making it a suitable data engine for magnetic models. As an example, we train a DeePSPIN model for the iron system and predict the Curie temperature through molecular dynamics simulations.

### v3

@@ Abstract
We propose a non-collinear spin-constrained method that generates training data for deep-learning-based magnetic model, which provides a powerful tool for studying complex magnetic phenomena that requires large-scale simulations at the atomic level. First, we propose a basis-independent projection method for calculating atomic magnetic moments by applying a radial truncation to numerical atomic orbitals. A double-loop Lagrange multiplier method is utilized to ensure the satisfaction of constraint conditions while achieving accurate magnetic torque. The method is implemented in ABACUS with both plane wave basis and numerical atomic orbital basis. We benchmark the iron (Fe) systems and analyze differences from calculations with the plane wave basis and numerical atomic orbitals basis in describing magnetic energy barriers. Based on an automated workflow composed of first-principles calculations, magnetic model, active learning, and dynamics simulation, more than 30,000 first-principles data with the information of magnetic torque are generated to train a deep-learning-based magnetic model DeePSPIN for the Fe system. By utilizing the model in large-scale molecular dynamics simulations, we successfully predict Curie temperatures of $\alpha$-Fe close to experimental values.
1 Introduction
@@
The study of magnetic materials has been a cornerstone of condensed matter physics, with implications ranging from fundamental science to technological applications. Density Functional Theory (DFT) [1964-hk, 1965-ks] has emerged as a powerful tool for understanding the electronic and magnetic properties of these materials [2023-misha]. However, accurate description of excited magnetic states within DFT has remained a formidable challenge due to the complex interplay between electron and lattice behaviors.
@@
The development of constrained Density Functional Theory (cDFT) has marked a significant advancement in this field  [1984-dederichs, 2012-constrained-review]. In the pioneering work of Dederichs et al. in 1984 [1984-dederichs], DFT was extended to arbitrary constraints through the introduction of the Lagrange multiplier method. Since then, several cDFT methods have been used to study excited states of charge distribution [1984-dederichs, 2005-wu] or magnetization [1998-stocks, 2004-kurz, 2018-SIESTA, 2000-gebauer, 2015-ma, 2020-SPHInX, 2023-cai]. Introducing penalty functions is commonly adopted to implement any constraint. [2000-gebauer, 2015-ma]. Wu and Van Voorhis (2005) [2005-wu] introduced an efficient algorithm, which allows for the use of a double-loop method to find the effective Lagrange multiplier that satisfies the constraints. In addition, the cDFT methods have been implemented based on different basis sets, such as the real-space grid [2000-gebauer], the full-potential linearized augmented plane-wave (FLAPW) [2004-kurz], the projector augmented-wave (PAW) [2015-ma, 2020-SPHInX, 2023-cai], and the numerical atomic orbitals (NAOs) [2018-SIESTA]. These efforts provide powerful tools in understanding charge and magnetization fluctuations in solids, predicting spin-dependent phenomena, and characterizing electron transfer reactions in molecules [2012-constrained-review, 1998-stocks, 2004-kurz, 2005-wu, 2015-ma, 2018-SIESTA, 2020-SPHInX, 2023-cai].
@@
In recent years, the integration of deep learning models with DFT has offered unprecedented opportunities for transferring first-principles accuracy to larger scales [2024-dpa1, 2023-dpa2, 2024-Qi, 2022-Chen, 2023-Deng, 2023-Batatia, 2023-Choudhary, 2021-choudhary]. The available databases [2013-mp, 2020-materials-cloud, atomly, 2024-OMat24] contain a large amount of DFT data that can be used to train machine learning models for different purposes. [2024-unisemi, wu2023universal, Wu23p144102]. Recently, some magnetic models [2024-fe-npj, 2024-deepspin] have emerged that incorporate the degrees of freedom of spin into the existing neural networks and can perform large-scale magnetic dynamics simulations. However, these models suffer from a lack of data since most existing DFT data in datasets are non-magnetic or only include ground-state magnetic results. Ground-state data cannot meet the sampling requirements of magnetic models with an additional degree of freedom spin. From this perspective, spin-cDFT methods not only act as powerful first-principles tools for studying magnetic excited states at the atomic scale but also provide training data for machine-learning-based magnetic models.
To address the substantial data requirements for AI-driven magnetic modeling, cDFT methodology must simultaneously satisfy multiple critical requirements: (1) robust convergence and high accuracy for reliable large-scale sampling, (2) consistent precision across datasets to eliminate untraceable errors, (3) minimal dependence on manually tuned parameters for automated active learning workflows, (4) adaptable constraint conditions accommodating diverse magnetic systems, and (5) optimized computational efficiency to address the inherent cost premium of magnetic versus non-magnetic calculations. To establish a robust data generation framework meeting these criteria, we have developed and implemented a Lagrange multiplier-based cDFT approach within the open-source ABACUS software [Chen2010, Li2016, lin_accuracy_2020, 2025-zhou], which provides energies, atomic forces, stresses, and magnetic torques for any magnetic excited state, making it a suitable data engine for magnetic models. Additionally, we propose a basis-independent local orbital projection method for calculating magnetic moment magnitudes. Unlike the Mulliken population method [2018-SIESTA], which, while maintaining the sum rule, may introduce contributions from non-local orbitals. Benefiting from the locality of NAOs, magnetic moments are obtained through the projection of atomic orbitals. By appropriately truncating the local atomic orbitals and smoothing near the truncation, we find that the modulated atomic orbitals provide satisfactory performance in stability and convergence during cDFT calculations. We present a unified implementation of cDFT that works efficiently with both plane-wave (PW) and numerical atomic orbital (NAO) basis. The spin-cDFT is fully optimized using the double-loop approach [2005-wu]. The optimized torque holds significant value for magnetic dynamics simulations and configuration space exploration. While conventional penalty methods [2015-ma] yield approximate torques $\lambda$, their fixed $\lambda$ induces non-negligible errors in energy derivatives. The adaptive scheme circumvents this by enforcing $\Delta M\to 0$. Moreover, the penalty methods encounter convergence difficulties when $\lambda$ is too large [2020-streib-prb], which requires careful handling of the iterative process. More critically, while the penalty function method can achieve reliable calculations for case studies through careful adjustment of $\lambda$, this system-dependent tuning process becomes impractical for the AI-driven large-scale sampling. A predetermined $\lambda$ exhibits varying accuracy levels across different configurations, introducing untraceable error that compromise subsequent model training. The limited accuracy, unstable convergence, and untraceable errors in large datasets render the penalty function method inadequate for meeting requirements 1/2/3. In contrast, the Lagrange multiplier method effectively addresses these challenges through its dual-optimization scheme [2004-kurz, 2005-wu, 2020-SPHInX, 2023-cai].
Our implementation supports two optional constraint modes: the first constrains both the magnitude and direction of magnetic moments, while the second constrains only the moment direction. Despite Gyorffy et al.’s assertion that longitudinal and transverse spin fluctuations exhibit temporal separation [1985-gyorffy-first], recent studies indicate that longitudinal spin fluctuations play a significant role in processes such as ferromagnetic-paramagnetic transitions [2007-ruban-prb] and phonon-magnon interactions [2018-wu-magnon]. In this work, we primarily focus on the application of the full-constraint algorithm, enabling the construction of magnetic potential energy surfaces through integration with recently developed magnetic models [2024-deepspin]. In coupled spin-lattice dynamics [2025-huang-prep], the contribution of longitudinal spin fluctuations manifests in the thermodynamic statistical properties of equilibrium states across different temperatures within the energy domain. The real-time evolution of longitudinal fluctuations in the temporal domain falls outside the scope of the present study.
Through the synergistic integration of cDFT calculations, magnetic model, active learning algorithms, and coupled spin-lattice dynamics [2025-huang-prep], we have developed a comprehensive end-to-end framework for autonomous magnetic model development. The workflow achieves complete automation throughout the entire training process, eliminating the need for manual intervention. Starting from scratch, we successfully trained two distinct models using PW and NAO basis, respectively. Both models yielded quantitatively consistent ferromagnetic-paramagnetic phase transition. This robust agreement across independent models provides validation of our methodology’s accuracy and transferability. This research paradigm offers valuable insights for AI-driven magnetic materials simulation. Furthermore, our work provides the research community with a reliable and accessible tool for related investigations.